OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (anonymous):

First find the slope.\[\frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\]

OpenStudy (anonymous):

The slope = m so substitute the answer from the previous formula to the next. \[y - y _{1} = m ( x - x _{1})\]

OpenStudy (anonymous):

You can use any points either (1,3) or (4,4) to substitute for\[y _{1} x_{1}\]